OpenStudy (anonymous):

Find the Solution of this Equation. 4x-1=-9

OpenStudy (anonymous):

add 1 to both sides....this will get 4x by itself. Then divide both sides by 4, getting x by itself. And that is what you want...x by itself....that will be your answer.

OpenStudy (anonymous):

Do you add it to the 4x And 1 ?

OpenStudy (anonymous):

you cannot add 4x + 1....but you can add -1 + 1. And don't forget to add the 1 to the other side

OpenStudy (anonymous):

let me rephrase that.... 4x - 1 + 1 = - 9 + 1 now, combine like terms

OpenStudy (anonymous):

4x-1=-9 +1 +1 4x= -8 4x/4= -8/4 x= -2

OpenStudy (anonymous):

I could have just given him the answer....I was trying to explain it.